RC Update and Run Plan for Aug 5, 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

Wednesday Aug 5 2020.

  • Hall C in beam permit waiting to re-establish beam.
  • MCC is recovering the machine.

General updates

  • Finishing 1-pass beam to Hall B and will establish 30 uA (planned) to Hall C.
  • Pass change for B before bringing A online. Priority will be beam at desired current to C.

Pending problem

  • Need beam. Want beam to 42 uA for BCM calibration.

Recent polarization measurements

  • ~34%

Current Run Plan

Normal Production running

  • Prerequisites
    • Make sure the raster size is correct (4-4.5)
    • Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
    • Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
    • Always take 120Hz of EDTM
    • Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
    • Change the IHWP every 4 hours when running on glass cells
  • Run plan
    • Re-establish beam [1].
    • When first able to send CW beam to dump, target in no target position, 2min. on 1uA, 1.5uA, 2uA. Simona here for Solid Cer data.
    • If MCC can deliver up to 42 muA we will do a BCM calibration [2]
    • Production on 3He; if not able to reach 30 uA, pro-rate the ABUs

Notes for tomorrow's 7:45 meeting

  •  ?


  • After ~ 10 days of running Hall C will go to 1 pass for ~3 days

Hall Access

  • None planned right now

When there is a down:

  • If the down is 2+ hours call Simona: she will check with the RC to see if there is an opportunity for work in the Hall (SoLID Cherenkov parasitic test)

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[3]

Planned Accelerator Configuration Changes

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links